OpenStudy (anonymous):

Use the Quadratic Formula to solve the equation. -2x^2 - 5x + 5 = 0

OpenStudy (anonymous):

@dtan5457

OpenStudy (anonymous):

@Nnesha

Nnesha (nnesha):

do you know the formula ??

OpenStudy (anonymous):

I did it and i got \[-\frac{ 5 }{ 4 }\pm \frac{ \sqrt{64} }{ 4 }\] is that right?

Nnesha (nnesha):

hmm try to solve which is under square root again

Nnesha (nnesha):

that is not 64 :(

Nnesha (nnesha):

got it ???